OpenStudy (anonymous):

What is the simplified form of the quantity y-squared plus 2y minus 8 over the quantity y-squared plus 7y plus 12 ?

OpenStudy (anonymous):

\[\frac{y^2+2y-8}{y^2+7y+12}\]We have to find the roots\[y^2+2y-8=0\]Roots\[y_1=-4~~and~~y_2=2\]And the second\[y^2+7y+12=0\]Roots\[y_3=-4~~and~~y_4=-3\]Now we can write this function\[\frac{(x+4)*(x-2)}{(x+4)*(x+3)}\]Simplifying\[\frac{(x-2)}{(x+3)}\]

OpenStudy (anonymous):

wow thanks!

OpenStudy (anonymous):

You're welcome ;)